solve the following equation
3/z-3 + 5/z-1 = 8/z-2
A) z=-4
B) z=-6
C) z=4
D) z=6

Answer by edjones(7311)

3/(z-3) + 5/(z-1) = 8/(z-2)
(3(z-1)(z-2))+(5(z-3)(z-2))=8(z-3)(z-1) Multiply each side by the LCM to eliminate fractions.
3(z^2-3z+2)+5(z^2-5z+6)=8(z^2-4z+3)
3z^2-9z+6+5z^2-25z+30=8z^2-32z+24
8z^2-34z+36=8z^2-32z+24
-2z+12=0 subtract the right side from the left.
-2z=-12
z=6
.
Ed